# An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks

### v1

@@ 4 Algorithm
In this section, we propose a main algorithm (Algorithm 1) to determine the equivalence of zero-one networks with two sub-algorithms (Algorithm 2 and 3). The goal of the main algorithm is to efficiently classify all $d$-dimensional $s$-species $m$-reaction zero-one networks according to the equivalence determined by their steady-state ideals. The correctness of Algorithm 1 follows from Theorem 3 and the correctness of the two sub-algorithms.
@@
The number of the species $s$, the number of the reactions $m$ and the dimension $d$
@@
Finitely many classes of $d$-dimensional $s$-species $m$-reaction zero-one networks such that the networks in the same class have the same steady-state ideal
@@
Enumerate all $d$-dimensional $s$-species $m$-reaction zero-one networks.
@@
Pick the networks satisfying the following two conditions simultaneously.
  1. 1.
    All the rows of the stoichiometric matrix $N$ are non-zero and change signs.
  2. 2.
    The set $\mathcal{F}^{+}(\mathcal{N})$ defined as in (10) is not empty.
(##\##By this step, we remove the networks definitely admitting no positive steady states.)
@@
For each remaining network, do the following steps:
Step 3.1.
If $c=m-1$ and $d\geq 2$, then delete the network.
Step 3.2.
If $c=m-2$ and $d=s\geq 3$, then delete the network.
(##\##Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
@@
We call CoefficientsForVacuous($s$, $d$), and we get a set $S\subset\mathbb{Q}^{s}$.
@@
For each remaining network, if there exists $a\in S$ such that
  • (1)
    $\mathcal{N}^{\top}a\neq 0$, and
  • (2)
    for any $i\in\{1,\dots,s\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the network.
@@
Delete the same networks after relabeling the species.
@@
We call CoefficientsForEquivalence($s$, $d$), and we get a set $S_{a}\subset\mathbb{Q}^{s}$.
Step 8.
For each pair of networks $G$ and $G^{^{\prime}}$ from the remaining networks in Step 6, do the following steps.
Step 8.1.
If there exists a permutation on the rows of $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ such that $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal after permutating their rows in the same way, then permutate the rows of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ in the same way, and go to Step 8.2. Otherwise, consider the next pair of networks.
Step 8.2.
Check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ satisfy the following condition: for any $i\in\{1,\dots,s\}$, there exist $(a_{i1}\dots,a_{is})^{\top}\in S_{a}$ and $(b_{i1},\dots,b_{is})^{\top}\in S_{a}$ such that
$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{s}b_{ij}row_{j}(\mathcal{N}^{^{\prime}}).$
If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. If not, then consider the next pair of networks.
Step 9.
For every class generated in Step 8, choose the first network as the representative element. Calculate the reduced Gröbner bases for these representative networks. Delete the networks whose reduced Gröbner base are $\{1\}$ or have a monomial as an element.
Step 10.
@@
A set $S\subset\mathbb{Q}^{s}$ such that for any $d$-dimensional $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that the conditions (1) and (2) of Theorem 1 hold, then there exists $\tilde{a}\in S$ satisfying the conditions (1) and (2) of Theorem 1
@@
Let $S:=\emptyset$. For every $n\in\{1,\dots,d-1\}$, choose $n$ linearly independent reactions from the reactions enumerated in Step 1. For every choice, do the following steps.
@@
Calculate the basis of the null space of $\mathcal{M}$ and add all the vectors in the basis into the set $S$.
(##\##Note here, all these vectors in the basis are in $\mathbb{Q}^{s}$.)
@@
Proof of the Correctness of Algorithm 2. Define
$\Delta_{1}$ :=$:=\{i\in\{1,\dots,m\}|col_{i}(\mathcal{Y})\neq\vec{0}\},$ (18)
$\Delta_{2}$ :=$:=\{1,\dots,m\}\setminus\Delta_{1}.$ (19)
Let $M$ and $M_{k}$ ($k\in\{1,2\}$) be the real vector spaces spanned by the columns of ${\mathcal{N}}$ and the set of column vectors $\{col_{i}({\mathcal{N}})|i\in\Delta_{k}\}$, respectively. By (18) and the condition (2) of Theorem 1, we have
@@
Since $\mathcal{N}^{\top}a\neq\vec{0}$ (the condition (1) of Theorem 1), we have $a\not\in M^{\perp}$. Since $M_{1}\subset M$, we have ${\tt dim}(M_{1})<{\tt dim}(M)=d$. So, by Step 2 of Algorithm 2, there exists $B\subset S$ such that $B$ is the basis of $\mathcal{M}^{\perp}_{1}$. Notice that there exists $\tilde{a}\in B$ such that $\tilde{a}\not\in M^{\perp}$, i.e., $\mathcal{N}^{\top}\tilde{a}\neq\vec{0}$. Otherwise, we have $B\subset M^{\perp}$ and hence, we have $M_{1}^{\perp}\subset M^{\perp}$, which is contrary to the fact that $a\in M_{1}^{\perp}\backslash M^{\perp}$. So, the conditions (1) and (2) of Theorem 1 hold for $\tilde{a}\in{\mathbb{Q}}^{s}$.
@@
    $\{0,1\}^{3}\cup(\{-1,1\}^{s}\times\{2\})\cup$
    ($(\{-1,1\}\times\{-2,2\}\times\{1\})\cup(\{-2,2\}\times\{-1,1\}\times\{1\})\cup$
@@
Algorithm 3.
CoefficientsForEquivalence
Input.
The number of species $s$ and the dimension $d$
Output.
A set $S_{a}\subset\mathbb{Q}^{s}$ such that if two $d$-dimensional $s$-species zero-one networks satisfy
  • (1)
    ${\mathcal{Y}}={\mathcal{Y}}^{^{\prime}}$, and
  • (2)
    for any $i\in\{1,\dots,s\}$,
    $row_{i}(\mathcal{N}^{^{\prime}})$ $\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$ (21)
then for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
$row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$
@@
Enumerate all possible zero-one reactions with $s$ species (here, we only enumerate one direction in a pair of reversible reactions). Write the corresponding stoichiometric vectors into a set $E\subset\mathbb{Q}^{s}$ (here, it is possible for different reactions to have the same stoichiometric vector, and we only put different vectors into the set $E$).
@@
Let $S_{a}:=\emptyset$. For any $d$ column vectors from $E$, do the following steps.
@@
Let $\mathcal{A}$ be the $s\times d$ matrix consisting of the $d$ column vectors. If $\mathcal{A}$ has any zero rows, then stop and go to the next group of $d$ vectors chosen from $E$.
@@
Denote by $\vec{1}$ and $-\vec{1}$ the column vectors whose coordinates are all $1$ and $-1$, respectively. Define
$\Lambda_{1}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=\vec{1}\},$ (22)
$\Lambda_{2}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=-\vec{1}\},$ (23)
$\Lambda_{3}$ :=$:=\{1,\dots,d\}\setminus(\Lambda_{1}\cup\Lambda_{2}).$ (24)
Define a set of vectors
$F:=\{0,1\}^{|\Lambda_{1}|}\times\{-1,0\}^{|\Lambda_{2}|}\times\{-1,0,1\}^{|\Lambda_{3}|}\subset\mathbb{Q}^{d}.$ (25)
(##\##Notice here $|\Lambda_{1}|+|\Lambda_{2}|+|\Lambda_{3}|=d$.)
For any column vector $b$ in the set $F$, do the following steps.
Step 2.2.1.
Solve $a\in\mathbb{Q}^{s}$ from $\mathcal{A}^{\top}a=b$.
Step 2.2.2.
If there exists finitely many solutions, then add the solutions into the set $S_{a}$. If there exists infinitely many solutions, then add one of the solutions into the set $S_{a}$. If there is no solutions, then go to the next vector $b\in F$.
@@
Return the set $S_{a}$.
Proof of the Correctness of Algorithm 3. Consider two $d$-dimensional $s$-species zero-one networks $G$ and $G^{^{\prime}}$. Assume that their reactant matrices are equal, i.e., $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$. First, we prove a claim.
Claim. If there exists $\mathcal{A}$ in Step 2.1 of Algorithm 3 such that $\mathcal{A}=\mathcal{N}$, then for any $j\in\{1,\dots,s\}$, there exists $b\in F$ (defined as in (25)) such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
In fact, by (22), for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{A})$ are $1$. Then, since $\mathcal{N}=\mathcal{A}$, for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{N})$ are $1$. By the definition of zero-one network, all the coordinates of $col_{i}(\mathcal{Y})$ are 00. Then, since $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, all the coordinates of $col_{i}(\mathcal{Y}^{^{\prime}})$ are 00. Thus, by the definition of zero-one network again, we have all the coordinates of $col_{i}(\mathcal{N}^{^{\prime}})$ are 00 or $1$. Hence, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{1}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is 00 or $1$. Similarly, we can derive that for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{2}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$ or 00. Note that by the definition of zero-one network, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{3}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$, 00 or $1$. Hence, by (25), for any $j\in\{1,\dots,s\}$, $row_{j}(\mathcal{N}^{^{\prime}})^{\top}\in F$. So, for any $j\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
Below, we prove the correctness of Algorithm 3 by discussing two cases.
  1. Case 1.
    Assume that all the columns of $\mathcal{N}$ are in the set $E$ generated in Step 1 of Algorithm 3.
@@
      Assume that all the columns of $\mathcal{N}$ are different. Since all the columns in $E$ are different. by the choice of $\mathcal{A}$ in Step 2 of Algorithm 3, there exists $\mathcal{A}$ such that $\mathcal{A}=\mathcal{N}$. By (21), for any $i\in\{1,\dots,s\}$, there exists $a^{^{\prime}}:=(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (26)
      Hence, by (26), for any $i\in\{1,\dots,s\}$, we have
      $\mathcal{N}^{\top}a^{^{\prime}}=row_{i}(\mathcal{N}^{^{\prime}})^{\top}.$ (27)
      By $\mathcal{A}=\mathcal{N}$ and by the claim, for any $i\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$. Hence, by $\mathcal{N}=\mathcal{A}$ and (27), we have $\mathcal{A}^{\top}a^{^{\prime}}=b$, i.e., there exists a solution to the equation $\mathcal{A}^{\top}a=b$. By Step 2.2.2, there exists $a\in S_{a}$ such that $\mathcal{A}^{\top}a=b$. Hence, for any $i\in\{1,\dots,s\}$, by $\mathcal{A}=\mathcal{N}$ and $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$, we have $\mathcal{N}^{\top}a=row_{i}(\mathcal{N}^{^{\prime}})^{\top}$, i.e., $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}=a\in S_{a}$.

### v2

@@ 4 Algorithm
In this section, we propose Algorithm 1, together with two sub-algorithms (Algorithm 2 and Algorithm 3), to determine the equivalence of zero-one networks. The goal is to efficiently classify all $d$-dimensional zero-one networks consisting of $s$ species $m$ reactions according to the equivalence determined by their steady-state ideals. The correctness of Algorithm 1 follows from Theorems 13, Lemma 1 and the correctness of the sub-algorithms.
@@
The number of species $s$, the number of reactions $m$ and the dimension $d$
@@
Finitely many classes of the steady-state systems $f$ corresponding to the $d$-dimensional $s$-species $m$-reaction zero-one networks such that the systems in the same class generate the same steady-state ideal
@@
Enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ for all $s$-species zero-one reactions and compute the corresponding stoichiometric vectors $\beta-\alpha$. Then, enumerate all possible $s\times m$ stoichiometric matrices $\mathcal{N}$ with rank $d$ whose columns are formed by these stoichiometric vectors.
@@
We call CoefficientsForInconsistent($s$, $d$, $m$), and we get a set $S_{1}\subset\mathbb{Q}^{s}$. For each stoichiometric matrix obtained in Step 1, we do the following steps.
Step 2.1.
If there exists $a\in S_{1}$ such that the vector $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then delete this matrix.
Step 2.2.
If the set $\mathcal{F}^{+}(\mathcal{N})$ is empty, then delete this matrix.
(#\#By Step 2, we remove the stoichiometric matrices whose all possible corresponding networks definitely admit no positive steady states by Theorem 1 and a standard way.)
@@
For each remaining stoichiometric matrix, delete the stoichiometric matrices that have the same form with it.
@@
For every remaining stoichiometric matrix $\mathcal{N}$, generate all possible corresponding reactant matrices $\mathcal{Y}$ for zero-one networks. For every pair $\{\mathcal{N},\mathcal{Y}\}$, we do the following steps.
Step 4.1.
If
  1. (I)
    c=m1c=m-1, and $d\geq 2$, or
  2. (II)
    c=m2c=m-2, and $d=s\geq 3$,
then delete the pair. (#\#Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
Step 4.2.
We call CoefficientsForVacuous($s$, $d$), and we get a set $S_{2}\subset\mathbb{Q}^{s}$. If there exists $a\in S_{2}$ such that
  1. (I)
    $\mathcal{N}^{\top}a\neq\vec{0}$, and
  2. (II)
    for any $i\in\{1,\dots,s\}$, the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero implies $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the pair.
(#\#By Step 4, we remove the networks that have vacuous steady-state ideals according to Theorems 23.)
@@
Combine the remaining pairs $\{\mathcal{N},\mathcal{Y}\}$ into the same group if they have the same reduced row echelon form of $\mathcal{N}$. For each group, we do the following steps.
Step 5.1.
For each pair $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$, we do the following steps.
Step 5.1.1.
Check if there exists a permutation on the rows of $\mathcal{Y}$ such that $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal after permuting the rows of $\mathcal{Y}$.
Step 5.1.1.1.
If yes, then put the pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$ into the same group.
Step 5.1.1.2.
If not, then consider the next pair of matrices.
(#\#By Step 5.1, we check the equivalence according to Lemma 1.)
Step 5.2.
For every group, choose the first pair as the representative element. Note that every pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ uniquely determines a network.
@@
For every representative network obtained from Step 5, calculate the reduced Gröbner bases for steady-state ideal $\langle f\rangle$. Delete the systems whose corresponding reduced Gröbner bases are $\{1\}$ or have a monomial as an element.
(#\#Note here, if the Gröbner basis contains a single monomial, then the network corresponding to the system admits no positive steady states for a generic choice of rate constants.)
@@
Remark 5.
Note that in Step 1, we first enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$, rather than the stoichiometric vectors directly. This approach facilitates extensions of the algorithm, particularly for the quadratic zero-one network where we require that the sum of $\alpha_{ij}$ for every reactant equals $2$.
Remark 6.
We remark that Step 2.1 in Algorithm 1 generalizes [16, Algorithm 1–Step 2.1], which follows from Remark 1 and the correctness of the following Algorithm 2.
Algorithm 2.
CoefficientsForInconsistent
Input.
The number of species $s$ and the number of the reactions $m$
Output.
A set $S\subset\mathbb{Q}^{s}$ such that for any $m$-reaction $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that the condition of Theorem 1 holds, then there exists $\tilde{a}\in S$ satisfying the condition of Theorem 1.
Step 1.
Enumerate all possible $s$-species zero-one stoichiometric vectors.
Step 2.
Let $S=\{e_{1},\dots,e_{s}\}$, where $e_{i}\in\mathbb{Q}^{s}$ satisfies the $i$-th coordinate is $1$ and other coordinates are 0. Let $w:=(w_{1},\ldots,w_{s})$ be an indeterminate vector in $\mathbb{Q}_{\geq 0}^{s}$. Enumerate all relations among $w_{1},\ldots,w_{s}$. For every relation, we do the following steps.
Step 2.1.
For each stoichiometric vector $\gamma$ generated in Step 1, if it satisfies $w^{\top}\gamma<0$, then delete it. If not, then keep it.
Step 2.2.
For the remaining stoichiometric vectors in Step 2.1, enumerate all possible choices of $m$ vectors from them, forming a stoichiometric matrix $\mathcal{N}$. For each $\mathcal{N}$, we do the following step.
Step 2.2.1.
If all non-zero rows of $\mathcal{N}$ change signs, then check if there exists $q\in\mathbb{Q}^{s}$ such that $q^{\top}\mathcal{N}$ is non-zero and does not change sign. (#\#Here, we use FindInstance in Mathematica.) If so, add the Cartesian product $\bigotimes\limits_{i=1}^{s}\{q_{i},-q_{i}\}$ to the set $S$.
Proof of the Correctness of Algorithm 2 For any $m$-reaction $s$-species zero-one network with the stoichiometric matrix $\mathcal{N}$, assume that there exists $a\in\mathbb{Q}^{s}$ such that the condition of Theorem 1 holds, i.e., $a^{\top}\mathcal{N}$ is non-zero and does not change sign. Note that if there exists $i\in\{1,\dots,s\}$ such that ${\rm row}_{i}({\mathcal{N}})$ is non-zero and does not change sign, then $e_{i}^{\top}\mathcal{N}={\rm row}_{i}({\mathcal{N}})$ is non-zero and does not change sign. Therefore, by Step 2, the vector $\tilde{a}=e_{i}\in S$ satisfies the condition of Theorem 1. Below, we assume that all non-zero rows of $\mathcal{N}$} change signs.
  • (I)
    Assume that all entries of $a^{\top}\mathcal{N}$ are non-negative. Let $\hat{a}=|a|\in\mathbb{Q}_{\geq 0}^{s}$. Then, there exists a relation of $w$ in Step 2 that satisfies the same relation of $\hat{a}$. Let $\hat{\mathcal{N}}$ be the matrix formed by the rows ${\rm sign}(a_{i}){\rm row}_{i}({\mathcal{N}})$ $(i=1,\ldots,s)$. Note that
    $a^{\top}\mathcal{N}=\hat{a}^{\top}\hat{\mathcal{N}}$.
    Since $a^{\top}\mathcal{N}$ is non-zero and all its entries are non-negative, the same holds for $\hat{a}^{\top}\hat{\mathcal{N}}$. Hence, for every column vector $\gamma$ of the matrix $\hat{\mathcal{N}}$, we have $\hat{a}^{\top}\gamma\geq 0$. Then, by Step 2.1 and Step 2.2, $\hat{\mathcal{N}}$ is one of the enumerated matrices under the relation of $\hat{a}$. Since all non-zero rows of $\mathcal{N}$ change signs, then every non-zero row of $\hat{\mathcal{N}}$ also changes sign. Then, by Step 2.2.1.2, the vector $q\in\mathbb{Q}^{S}$ obtained via FindInstance satisfies that $q^{\top}\hat{\mathcal{N}}$ is non-zero and all entries are non-negative. Let $\tilde{a}=({\rm sign}(a_{1})q_{1},\ldots,{\rm sign}(a_{s})q_{s})$. By Step 2.2.1.2, we know that $\tilde{a}\in S$. Notice that $\tilde{a}^{\top}\mathcal{N}=q^{\top}\hat{\mathcal{N}}$. Hence, $\tilde{a}^{\top}\mathcal{N}$ is non-zero and all its entries are non-negative, i.e., $\tilde{a}$ satisfies the condition of Theorem 1.
  • (II)
    Assume that all entries of $a^{\top}\mathcal{N}$ are non-positive. Notice that $a^{\top}\mathcal{N}$ is non-zero and all entries of $a^{\top}\mathcal{N}$ are non-negative. Then, the proof is similar to (I).
Example 5.
We illustrate how Algorithm 2 works for $m=5$ and $s=3$.
Step 1.
Enumerate all possible $3$-species stoichiometric vectors according to zero-one reactions, i.e., all vectors in
$\{-1,0,1\}^{3}\setminus(0,0,0)^{\top}.$
Step 2.
Let $S=\{(0,0,1)^{\top},(0,1,0)^{\top},(0,0,1)^{\top}\}$. Enumerate all relations among $w_{1}$, $w_{2}$, and $w_{3}$, where $w_{i}\in\mathbb{Q}_{\geq 0}$. By symmetry, we only need to consider the following cases.
$w_{1}>w_{2}>w_{3},\ w_{1}=w_{2}>w_{3},\ w_{1}>w_{2}=w_{3},\ w_{1}=w_{2}=w_{3}.$
We treat every relation as follows.
  • (I)
    $w_{1}>w_{2}>w_{3}$.
    Step 2.1.
    For each stoichiometric vector $\gamma$ generated in Step 1, if it satisfies $w^{\top}\gamma<0$, then delete it. If not, then keep it. For example,
    • (i)
      we delete $(-1,-1,-1)^{\top}$ since $w^{\top}(-1,-1,-1)^{\top}=-w_{1}-w_{2}-w_{3}<0$;
    • (ii)
      we keep $(-1,1,1)^{\top}$ since we cannot determine the sign of $w^{\top}(-1,1,1)^{\top}=-w_{1}+w_{2}+w_{3}$ under the condition $w_{1}>w_{2}>w_{3}\geq 0$.
    The remaining vectors are
    $(-1,1,1)^{\top},$ $(0,0,1)^{\top},$ $(0,1,-1)^{\top},$ $(0,1,0)^{\top},$ $(0,1,1)^{\top},$
    $(1,-1,-1)^{\top},$ $(1,-1,0)^{\top},$ $(1,-1,1)^{\top},$ $(1,0,-1)^{\top},$ $(1,0,0)^{\top},$
    $(1,0,1)^{\top},$ $(1,1,-1)^{\top},$ $(1,1,0)^{\top},$ $(1,1,1)^{\top}.$
    Step 2.2.
    For the remaining stoichiometric vectors in Step 2.1, enumerate all possible choices of $5$ vectors from them, forming a stoichiometric matrix $\mathcal{N}$. For each $\mathcal{N}$, we do Step 2.2.1 of Algorithm 2. For example, consider
    $\mathcal{N}=\begin{pmatrix}-1&0&0&0&1\\ 1&0&1&1&-1\\ 1&1&-1&0&-1\end{pmatrix}$.
    Step 2.2.1.
    Notice that all non-zero rows of $\mathcal{N}$ change signs. Using FindInstance in Mathemetica, we find $q=(2,1,1)^{\top}$ such that $q^{\top}\mathcal{N}$ is non-zero and does not change sign. Then, add the Cartesian product $\{2,-2\},\times\{1,-1\},\times\{1,-1\}$ to set $S$.
    After enumerating all possible matrices, we add the follow set of vectors into $S$:
    $\{0,\pm\frac{1}{4},\pm\frac{1}{3},\pm\frac{1}{2},\pm\frac{2}{3},\pm\frac{3}{4},\pm 1,\pm\frac{3}{2},\pm 2,\pm 3\}^{3}.$
  • (II)
    We implement the similar steps as case (I) for the other relations $w_{1}=w_{2}>w_{3}$, $w_{1}>w_{2}=w_{3}$ and $w_{1}=w_{2}=w_{3}$. And finally, we get
    $S=\{0,\pm\frac{1}{4},\pm\frac{1}{3},\pm\frac{1}{2},\pm\frac{2}{3},\pm\frac{3}{4},\pm 1,\pm\frac{3}{2},\pm 2,\pm 3\}^{3}.$ (19)
@@
A set $S\subset\mathbb{Q}^{s}$ such that for any $d$-dimensional $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that conditions (I) and (II) of Theorem 2 hold, then there exists $\tilde{a}\in S$ satisfying conditions (I) and (II) of Theorem 2
@@
Let $S:=\emptyset$. For every $n\in\{1,\dots,d-1\}$, for all possible choices of $n$ linearly independent reactions from the reactions enumerated in Step 1, we do the following steps. Note that here we operate on all possible combinations of $n$ reactions.
@@
Calculate a basis of the nullspace of $\mathcal{M}$ in $\mathbb{Q}^{s}$ and add all the vectors in the basis into the set $S$.
@@
Proof of the Correctness of Algorithm 3. For any $d$-dimensional $s$-species zero-one network, suppose $\mathcal{Y}$} and $\mathcal{N}$} are the reactant matrix and the stoichiometric matrix. Define
$\Delta_{1}$ $:=\{i\in\{1,\dots,m\}|col_{i}(\mathcal{Y})\neq\vec{0}\},$ (20)
$\Delta_{2}$ $:=\{1,\dots,m\}\setminus\Delta_{1}.$ (21)
Let $M$ and $M_{k}$ ($k\in\{1,2\}$) be the real vector spaces spanned by the columns of $\mathcal{N}$} and the set of column vectors $\{col_{i}({\mathcal{N}})|i\in\Delta_{k}\}$, respectively. If there exists $a\in\mathbb{Q}^{s}$ such that conditions (I) and (II) of Theorem 2 hold, then by (20) and condition (II) of Theorem 2, we have
@@
Since $\mathcal{N}^{\top}a\neq\vec{0}$ (condition (I) of Theorem 2), we have $a\not\in M^{\perp}$. Since $M_{1}\subset M$, we have ${{\rm dim}}(M_{1})<{{\rm dim}}(M)=d$. So, by Step 2 of Algorithm 3, there exists $B\subset S$ such that $B$ is a basis of $M^{\perp}_{1}$. Notice that there exists $\tilde{a}\in B$ such that $\tilde{a}\not\in M^{\perp}$, i.e., $\mathcal{N}^{\top}\tilde{a}\neq\vec{0}$. Otherwise, we have $B\subset M^{\perp}$ and hence, we have $M_{1}^{\perp}\subset M^{\perp}$, which is contrary to the fact that $a\in M_{1}^{\perp}\backslash M^{\perp}$. So, conditions (I) and (II) of Theorem 2 hold for $\tilde{a}\in{\mathbb{Q}}^{s}$.
@@
$\{0,1\}^{3}\cup(\{-1,1\}^{{2}}\times\{2\})\cup$
$(\{-1,1\}\times\{-2,2\}\times\{1\})\cup(\{-2,2\}\times\{-1,1\}\times\{1\})\cup$
@@
5 Implementation
We have implemented the procedure outlined in Algorithm 1, with supporting code publicly available at https://github.com/YueJ13/equivalence. We further apply this algorithm to all three-dimensional zero-one networks consisting of three species and five reactions.
@@ 5 Implementation
Enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ for all $3$-species zero-one reactions and compute the corresponding stoichiometric vectors $\beta-\alpha$. Then enumerate all possible $3\times 5$ stoichiometric matrices $\mathcal{N}$ of rank $3$ whose columns are stoichiometric vectors. There are $43520$ matrices.
@@
We call CoefficientsForInconsistent($3$, 33, 55), and by the computation carried out in Example 5, we get a set of rational vectors $S_{1}$ in (19). For each stoichiometric matrix obtained in Step 1, we do the following steps.
@@
If there exists $a\in S_{1}$ such that the vector $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then delete this matrix.
@@
If the set $\mathcal{F}^{+}(\mathcal{N})$ is empty, then delete this matrix.
Then, we get $14444$ matrices.
@@
For each remaining stoichiometric matrix, delete the stoichiometric matrices that have the same form with it. Then, we get $2548$ matrices.
Step 4.
For every remaining stoichiometric matrix $\mathcal{N}$, generate all possible corresponding reactant matrices $\mathcal{Y}$. For every pair $\{\mathcal{N},\mathcal{Y}\}$, we do the following steps.
Step 4.1.
If $c=3$ or $4$, then delete the pair. (#\#Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
Step 4.2.
We call CoefficientsForVacuous($3$, 33), and by the compotation carried out in Example 6, we get a set of rational vectors $S_{2}$ in (Step 2.). If there exists $a\in S_{2}$ such that
@@
    $\mathcal{N}^{\top}a\neq\vec{0}$, and